In the table below calculate opportunity cost for: (12) Country Computers Sewing Machines Japan 500 900...

  1. In the table below calculate opportunity cost for: (12)
Country Computers Sewing Machines
Japan 500 900
Canada 300 480

a. Japan of producing a computer rather than a sewing machine

b. Canada of producing a computer rather than a sewing machine

c. Japan of producing a sewing machine rather than a computer

d. Canada of producing a sewing machine rather than a computer

e. Which country has comparative advantage in producing sewing machines?

f. Which country has comparative advantage in producing computers?
Answer #1

Opportunity cost is the cost of the benefit foregone.

a) Opportunity cost of Japan producing a computer rather than a sewing machine

Opportunity cost = sewing machines given up ÷ computers produced

= 900÷ 500 = 1.8 sewing machines

b) Opportunity cost of Canada producing a computer rather than a sewing machine

Opportunity cost = sewing machines given up ÷ computers produced

= 480 ÷ 300 = 1.6 sewing machines

c) Opportunity cost of Japan producing a sewing machine rather than a computer

Opportunity cost = computers given up ÷ sewing machines produced

= 500 ÷ 900 = 0.56 computers

d) Opportunity cost of Canada producing a sewing machine rather than a computer

Opportunity cost = computers given up ÷ sewing machines produced

= 300 ÷ 480 = 0.625 computers

e) Since the opportunity cost of producing a sewing machine is lower for Japan when compared to Canada, Japan has a comparative advantage in producing sewing machines

f) Canada has a comparative advantage over Japan while producing computers because of lower opportunity cost of producing computer when compared to Japan.

Please provide feedback if it was helpful. Feel free to comment if you have any doubts.